MINUTES OF THE PS-ISOLDE
COORDINATION MEETING held on 8
Butler, R. Catherall, B. Crepieux, A. Dorsival, V. Fedosseev,U. Georg, U.
Koester, G. Le Scornet, M. Lindroos, S. Meunier, T. Nilsson, J. Parra-Lopez, F. Wenander
ISOLDE at URL: http://isolde.cern.ch
Minutes at URL: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html
Consolidation projects at URL: http://isolde-consolidation.web.cern.ch/isolde-consolidation/
PSMB meetings at URL: http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html
The HRS did run with a UC target for the production of K
for an experiment using the Tonnere neutron detector. The run was successful
but had to be terminated early as the K yield deteriorated sharply on
Saturday morning. This was after a long run with a lot of protons on an very
hot target and line. Several collections for solid state physics could be
performed during the run thanks to the flexibility shown by the primary user. An
autopsy will be done in due time on this target unit to establish the reason
for the yield deterioration
The magnet program on the HRS blocked during the run.
It seems prudent to invest some of the available consolidation money for the
HRS to update the Metrolab NMR-probe hardware. Tim will investigate this.
Ta foil target was mounted on the GPS. Problems were experienced again with
the automatic heating program. This program will be replaced in the shutdown
by a JAVA version running on the linux consoles.
- The new control software for the tape station is now being tested. It is
possible to improve the electronics for the event counter to cope with very
high rates, something that could be of great interest for some applications.
4. Tape station
- The tape station was made operational again on Monday. The settings for the
hardware were totally deregulated even though no changes had been made to
the control potentiometers on the hardware. The hardware is in strong need
of major consolidation. Mats and Gerit Jan promised to investigate what possibilities exist
in the short term future to improve the situation.
- A power supply broke down and had to be replaced during the setting-up for
the Be run.
6. Off-line facilities
- Some of the new power supplies have now been specified to cope with
current higher than 5 mA (e.g. to cope with the Mono ECR source).
- You can submit abstracts
on the SAFERIB web site (http://cern.ch/saferib). The program will be made available ASAP on the web site.
- A warm welcome to Ivan Podadera Aliseda who has joined the ISOLDE section in
the operation group. Ivan, who is from Barcelona, will do his PhD with
Ari on the design and construction of the RFQ cooler.
- Remember to contact the vacuum group for ALL none-routine interventions on
the ISOLDE vacuum system. It might not be necessary for them to participate
in each intervention but they should always be informed.
- The magnet tests in the porch to the ISOLDE hall has been causing major disturbances
to the access for the hall. These particular magnet tests were agreed with
the responsible at ISOLDE in advance but, still, it serves as a reminder to all of
us to avoid such tests in this area in the future.
- The cleaning service for the hall are "improving". A new company
has been contracted by CERN for this task and they are now receiving
instructions for how they should intervene in the hall.
- A leak on the GPS front end faraday cup is being investigated. Jacques is organizing
a special meeting to discuss this issue.
15 October 2002 at 11.00
in the EP Conference Room A, building 304-1-001.
by Mats Lindroos